Sol Summary – February 15th

Crew IKEA Sol Summary 15 Feb 2018

SOL 5:

Summary Title: Shelter from the Storm

Author’s name: Atila Meszaros

Mission Status: All nominal

Sol Activity Summary:

Another slow and very windy day. The ladders installations and the mud have kept us inside most of the day. However, we managed to go out and we drive Opportunity, Spirit, and Deimos. Due to the new ladder installation, I’ve been trapped in the upper deck almost all day and we had to postpone our adventures until (I hope) tomorrow. All the Hab is screaming because of the winds. It sounds pretty awesome. but I hope I can get some sleep

Thank you very much for being with us tonight!


Look Ahead Plan:
Getting a new battery for Shannon’s car, taking Rovers out and I’m rooting for finally some field work and exploration!

Anomalies in work:
Installation of brand new ladders for the lower deck

Weather: Temperature: 5°C

Crew Physical Status: Healthy

Reports to be filed:

Sol Summary
Operations Report

Support Requested: None

Operations Report – February 15th

Crew IKEA Operations Report 15 Feb 2018

SOL: 5

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control
Notes on non-nominal systems: Nothing new to report

Generator (hours run): Generator was once again run all day

Solar— not checked % SOC
Diesel – 72 %
Propane – 47 psi.
Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.
Water (static) – 425 gallons
Trailer to Static Pump used – no
Water (loft) – Static to Loft Pump used – yes
Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: no
Oil Added? n/a
ATV Fuel Used: n/a Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report

Deimos rover used: Yes
Beginning charge: 100
Ending charge: 96
Currently charging: Yes

Sojourner rover used: No
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: Yes
Hours: 20.2
Beginning charge: 100
Ending charge: 75
Currently charging: No

Opportunity rover used: Yes
Hours: 14.9
Beginning charge:
Ending charge:
Currently charging: Yes

Curiosity rover used: No
Beginning charge:
Ending charge:
Currently charging: No

HabCar used and why, where? Yes, to town for packages. New jumper cables arrived for HabCar. Other set is in Hab.

General notes and comments: Septic tank is almost full. Pumping it out is scheduled for Monday. New battery charger (to replace the one that arrived broken) arrived today and will be stored in the RAM.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: The front airlock door was left unsecured while the contractors were working and the window got blown out again. We will fix it before the crew arrives.

The new stairs are taking shape! Both of the ladders have been installed on the upper deck and the base of the new stairs is in place but not installed. The ladder sections and the lower stairs still need to be installed as well as the railing at the top of the stairs. The contractor expects the installation to be complete on Saturday but the polishing/painting will be done later.

I should be able to send photos tomorrow.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Electrical connection has been scheduled for 3-6 weeks out.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Copyright © The Mars Society. All rights reserved. | Main Site